The following information was drawn from the year-end balance sheets of Finch River, Inc.:

Account Title 2017 2016
Bonds payable $ 735,000 $ 975,000
Common stock 201,000 122,000
Treasury stock 27,500 10,500
Retained earnings 87,900 61,500

Additional information regarding transactions occurring during 2017:

  1. Finch River, Inc., issued $47,300 of bonds during 2017. The bonds were issued at face value. All bonds retired were retired at face value.

  2. Common stock did not have a par value.

  3. Finch River, Inc., uses the cost method to account for treasury stock.

  4. The amount of net income shown on the 2017 income statement was $35,600.

Required

  1. Determine the amount of cash flow for the retirement of bonds that should appear on the 2017 statement of cash flows.

  2. Determine the amount of cash flow from the issue of common stock that should appear on the 2017 statement of cash flows.

  3. Determine the amount of cash flow for the purchase of treasury stock that should appear on the 2017 statement of cash flows.

  4. Determine the amount of cash flow for the payment of dividends that should appear on the 2017 statement of cash flows.

  5. Prepare the financing activities section of the 2017 statement of cash flows.

Determine the amount of cash flow for the retirement of bonds, for the issue of common stock, for the purchase of treasury stock and for the payment of dividends that should appear on the 2017 statement of cash flows.

a. Cash flow for the retirement of bonds
b. Cash flow from the issue of common stock
c. Cash flow for the purchase of treasury stock
d. Cash flow for the payment of dividends

Prepare the financing activities section of the 2017 statement of cash flows. (Cash outflows should be indicated with minus sign.)

Cash Flows from financing activities
Net cash flow from investing activities
Answer #1

Bonds retired = Bonds payable, beginning + Bonds issued - Bonds payable, ending

= 975,000 + 47,300 - 735,000

= $287,300

Cash flow from the issue of common stock = Common stock, ending - Common stock, beginning

= 201,000 - 122,000

= $79,000

Cash flow for the purchase of treasury stock = Treasury stock, ending - Treasury stock, beginning

= 27,500 - 10,500

= $17,000

Cash payment for dividend = Retained earnings, beginning + Net income - Retained earnings, ending

= 61,500 + 35,600 - 87,900

= $9,200

a. Cash flow for the retirement of bonds - 287,300
b. Cash flow from the issue of common stock 79,000
c. Cash flow for the purchase of treasury stock - 17,000
d. Cash flow for the payment of dividends - 9,200
Cash Flows from financing activities
Cash flow for the retirement of bonds - 287,300
Cash flow from the issue of common stock 79,000
Cash flow for the purchase of treasury stock - 17,000
Cash flow for the payment of dividends - 9,200
Cash flow from issue of bonds 47,300
Net cash flow from financing activities - $187,200

Kindly comment if you need further assistance. Thanks